OpenStudy (parth1296):

Suppose you measure three numbers as follows: x= 200+-2 y= 50+-2 z=40+-2 where the three uncertainties are independent and random. Use step by step propagation to find the quantity q=x/(y-z) with its uncertainty.

OpenStudy (jiteshmeghwal9):

\[\frac{∆q}{q}=\frac{∆x}{x} + \frac{∆(y-z)}{(y-z)}\] Because during multiplication or division relative errors get added